A 20 kg brick stands upright on the ground as shown. What is the pressure it exerts on the ground? 10cm x 20cm x 5cm

1
Expert's answer
2022-04-12T11:30:22-0400

The pressure exerted by a brick on the ground

"p=\frac{F}{A}=\frac{mg}{A}"

"p=\frac{20*9.8}{0.2*0.05}=19600\:\rm Pa"